On land measuring 1282 sqft,, 940 sqft flat is built on each G+6 floor, Thus total built area is 940x7 = 6580 sqft. Out of this 400 sqft used as Commercial & 6180 as Residential.

question ---Out of 1282 Sqft land, how much will attributed as commercial & how much as residential?
1
Expert's answer
2020-03-06T15:47:59-0500

1) 400 sqft used as Commercial is 400/6580 = 0.061 or 6.1% of the total area,

2) 6180 sqft used as Residential is 6180/6580 = 0.939 or 93.9%.

3) So, out of 1282 Sqft land, 1282×0.061 = 78.2 Sqft will be attributed as commercial and 1282×0.939 = 1203.8 Sqft as residential.